# Integrity-protecting block cipher modes — Untangling a tangled web

### v1

@@ 1 Introduction
In either case we refer to the block added to the end of the plaintext as a Manipulation Detection Code (MDC). Whichever approach is adopted, the method for computing the MDC needs to be simple, or it offers no advantage over the more conventional ‘encrypt then MAC’ approach.
In all the modes we examine, the MDV (also known as an Integrity Control Value (ICV)) is defined to be a fixed, possibly secret, final plaintext block.
  • The first of the three schemes we examine is known as PES-PCBC. The name derives from it being a version of PCBC mode designed specifically for use in a Privacy Enhanced Socket (PES) protocol. There are actually a number of modes known as PCBC (Plaintext-Ciphertext Block Chaining); the version on which PES-PCBC was based is that incorporated in Kerberos version 4. This mode was shown to be insecure for the purposes of integrity protection by Kohl, [3]. For a discussion of the weaknesses of other variants of PCBC see [7]. The design goal for PES-PCBC was to enhance the security of PCBC by preventing the known attacks. This scheme and its properties are discussed in Section 3.
@@
In Section 6 we examine other more general attacks which apply to all, or at least large classes of, schemes sharing the same underlying structure as PES-PCBC, IOBC and EPBC. In doing so we establish that all three modes suffer from attacks which mean that they should not be adopted. We conclude the paper in Sections 7 and 8 by first briefly mentioning two other related modes (and their analyses), and then summarising the main conclusions that can be drawn from the analyses given.
@@ 3.1 Cryptanalysis
Suppose the ciphertext $C_{1},C_{2},\ldots,C_{t}$ was constructed using PES-PCBC from the plaintext $P_{1},P_{2},\ldots,P_{t}$, and that $j$ satisfies $1<j<t$. Suppose the $t+2$-block ciphertext $C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{t+2}$ is constructed as follows:
@@
    Hence $G^{\prime}_{j+1}=d_{K}(F^{\prime}_{j+1})=d_{K}(F_{j-1})=G_{j-1}$. Finally, we have $P^{\prime}_{j+1}=G^{\prime}_{j+1}\oplus F^{\prime}_{j}=G_{j-1}\oplus F_{j}=C_{j}$. Since $C^{\prime}_{j+2}=C_{j}$, $F^{\prime}_{j+1}=F_{j-1}$ and $G^{\prime}_{j+1}=G_{j-1}$, it is immediate that $F^{\prime}_{j+2}=F_{j}$, $G^{\prime}_{j+2}=G_{j}$, and $P^{\prime}_{j+2}=P_{j}$, and the desired result follows.
@@ 3.2 Impact
The above attack shows that, given just one PES-PCBC-encrypted message and knowledge of only a single plaintext block for this encrypted message, a ‘fake’ message can be constructed will be guaranteed to pass the integrity checks. This fact meant that it has been recognised since 1996/97 that PES-PCBC should not be used.
@@ 4.3 Cryptanalysis
Lemma 4.1 suggests how it may be possible to forge an IOBC-encrypted message so that the final block will contain the correct MDC. However, the problem remains of discovering $g(G^{\prime}_{u-1})\oplus g(G_{v-1})$, as used in constructing the forged ciphertext in the statement of the lemma. Recacha [10] discusses this very point, and explains that making this difficult motivates the inclusion of the function $g$ in the design of IOBC — that is, if $g$ was not included (as is the case for PES-PCBC), then simple forgeries could be achieved.
@@
It is not hard to see that if $g^{k}(G_{j})=G_{j}$ for some $k$, then Lemma 4.2(ii) could be combined with Lemma 4.1 to yield a forgery attack (given a ciphertext message with corresponding known plaintext). This point is made by Recacha [10], who explains that the bit permutation $g$ has been chosen so that the smallest integer $i>1$ such that $g^{i}$ is the identity permutation is $(n/2-1)(n/2+1)=n^{2}/4-1$ (since $m=n/2$ is even). The restriction that the maximum length of messages encrypted using IOBC is $n^{2}/2-1$, as defined in Section 4.2, prevents this problem arising in practice.
@@
The above result can now be used in a straightforward way to enable message forgeries. As described in detail in [9], if $n=64$, given an IOBC ciphertext containing at least 685 blocks and some of the corresponding plaintext, then a forged ciphertext can be created which will pass integrity checks with probability $2^{-22}$. 685 is clearly much less than the defined message length limit of 2047 blocks. A precisely analogous attack works for $n=128$, although the success probability is only $2^{-42}$.
@@ 5.2 A flawed cryptanalysis
The above Lemma implies that output bit pairs $(\ell^{\prime}_{i},r^{\prime}_{i})$ can never be equal to (0,1). In fact, we can obtain the following more general result which gives Lemma 5.1 as a special case.
Lemma 5.2 (Mitchell, [8])
@@
However, more recently, Di et al. [2] pointed out that the above analysis has a major flaw. The issue arises in the argument that, since $G_{j+1}=P_{j+1}\oplus F_{j}$, information about forbidden bit pairs in $F_{j}$, combined with knowledge of $P_{j+1}$, gives information about forbidden bit pairs in $G_{j+1}$. Di et al. [2] point out that if, there are two possibilities for a bit pair in $F_{j}$ then there will always still be two possibilities for the corresponding bit pair in $g(G_{j+1})$ — as opposed to the analysis in [8] which suggests that the number of possibilities will be reduced to one with probability 1/6. That is, the number of possibilities for a bit pair in $g(G_{j+1})$ will never go below two, preventing the attack strategy working.
@@ 6.1 A chosen plaintext forgery attack
    Under reasonable assumptions about the random behaviour of the encryption function, this will occur with probability $2^{-n}$. Hence, as $\Delta$ ranges over its $2^{n}$ possible values, the expected number of times that $C^{\prime}_{j}=C_{j}$ will hold is approximately 2, one of which will occur when $F^{\prime}_{j}=F_{i}$. The result follows.
We can now give the following simple result that uses the same notation as Lemma 6.1. Note that it uses Lemma 4.1, which we already observed holds regardless of the choice of $g$.
Lemma 6.2
Suppose that $C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{t^{\prime}}$ and $C_{1},C_{2},\ldots,C_{t}$ are as in the statement of Lemma 6.1, and suppose also that $P^{\prime}_{j}=P_{i}$ and $C^{\prime}_{j}=C_{i}$ for some $j<t^{\prime}$ and $i<t$. Then, with probability approximately 0.5, the constructed ciphertext message
@@
    The result follows immediately from Lemma 4.1, putting $u=j$, $v=i$ and observing that:
@@
    which equals $C_{i}$ with probability approximately 0.5, since, by Lemma 6.1 we know that $G^{\prime}_{j-1}=G_{i-1}$ with probability approximately 0.5.
@@ 6.2 A new attack approach with implications for EPBC
Suppose the ciphertext $C_{1},C_{2},\ldots,C_{t}$ was constructed using a scheme of the type shown in Figure 2, and that $P_{j}$ is a plaintext block for some $j$ satisfying $1<j<t$. Suppose the $t+2$-block ciphertext $C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{t+2}$ is constructed as follows:
@@
    although the precise value of $P^{\prime}_{j+1}$ is unimportant. The result follows trivially.
@@
Lemma 6.5
Suppose $g$ is as defined for EPBC, and (using the notation of Lemma 5.1) suppose also that $g(L||R)=L^{\prime}||R^{\prime}$. Then:
@@
    Claim (i) follows immediately, and claim (ii) follows from observing that $\lnot\ell\land r=1$ if and only if $r=1$ and $\ell=0$.
From Theorem 6.3, we can construct a possible forgery by guessing the value of $G_{j}\oplus g(G_{j})$. If $g$ is as defined for EPBC, then, from Lemma 6.5(i), we simply need to guess the first $n/2$ bits of $G_{j}\oplus g(G_{j})$. Moreover, if we restrict our guesses for this ‘first half’ to $n/2$-bit strings containing at most $n/8$ ones (where we assume that $n$ is a multiple of 8), then from Lemma 6.5(ii) we will have a better than evens chance of making a correct guess. The number of such strings is simply:
@@ 6.3 Issues with the use of Initialisation Vectors (IVs)
We can state the following result, which is essentially a special case of Lemma 4.1, and applies to all modes adhering to the design of Figure 2 and for which $g$ is linear (as is the case for IOBC and, trivially, PES-PCBC). In such a case, as observed in [9], the distributivity property $g(X\oplus Y)=g(X)\oplus g(Y)$ for any $X$ and $Y$ holds.
Lemma 6.7
@@
    The result follows from Lemma 4.1, setting $u=v=3$.
Thus if two ciphertexts are encrypted using the same IV, and a single plaintext block is known for each message, then a forgery can be constructed. Observe that this attack can be extended using Lemma 4.2ii.
@@ 7 Other related modes
There was a gap of some 16 years before the first of these two additional schemes was made public — IOC (short for Input and Output Chaining) was made public by Recacha in 2013 [11]. IOC is clearly closely related to IOBC and EPBC, but was designed to avoid the known issues with these schemes. IOC was made public at a time when there was a renewed interest in the area, at least partly prompted by a NIST initiative on lightweight cryptography [^2] As discussed at https://csrc.nist.gov/Projects/Lightweight-Cryptography, NIST began investigating cryptography for constrained environments in 2013, and one of the goals was to find lightweight methods for authenticated encryption.. A slightly revised version was made public in early 2p14, [14]. Cryptanalyses of both versions of the scheme first appeared in 2014 [1].

### v2

@@ Abstract
Keywords:
authenticated encryption cryptanalysis mode of operation
@@ 1 Introduction
In either case we refer to the block added to the end of the plaintext as a Manipulation Detection Code (MDC). Whichever approach is adopted, the method for computing the MDC needs to be simple, or it offers no advantage over the more conventional ‘encrypt then MAC’ approach. In all the modes we examine, the MDV (also known as an Integrity Control Value (ICV)) is defined to be a fixed, possibly secret, final plaintext block.
The first of the three schemes we examine is known as PES-PCBC. The name derives from it being a version of PCBC mode designed specifically for use in a Privacy Enhanced Socket (PES) protocol. There are actually a number of modes known as PCBC (Plaintext-Ciphertext Block Chaining); the version on which PES-PCBC was based is the one incorporated in Kerberos version 4. This mode was shown to be insecure for the purposes of integrity protection by Kohl, [3]. For a discussion of the weaknesses of other variants of PCBC see [7]. The design goal for PES-PCBC was to enhance the security of PCBC by preventing the known attacks. This scheme and its properties are discussed in Section 3.
@@
In Section 6 we examine other more general attacks which apply to all, or at least large classes of, schemes sharing the same underlying structure as PES-PCBC, IOBC and EPBC. In doing so we establish that all three modes suffer from attacks, and so they should not be adopted. We conclude the paper in Sections 7 and 8 by first briefly mentioning two other related modes (and their analyses), and then summarising the main conclusions that can be drawn from the analyses given.
@@ 3.2 Cryptanalysis
Suppose the ciphertext $C_{1},C_{2},\ldots,C_{t}$ was constructed using PES-PCBC from the plaintext $P_{1},P_{2},\ldots,P_{t}$, and that $j$ satisfies $1<j<t$. Suppose the $(t+2)$-block ciphertext $C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{t+2}$ is constructed as follows:
@@
Hence $G^{\prime}_{j+1}=d_{K}(F^{\prime}_{j+1})=d_{K}(F_{j-1})=G_{j-1}$. Finally, we have $P^{\prime}_{j+1}=G^{\prime}_{j+1}\oplus F^{\prime}_{j}=G_{j-1}\oplus F_{j}=C_{j}$. Since $C^{\prime}_{j+2}=C_{j}$, $F^{\prime}_{j+1}=F_{j-1}$ and $G^{\prime}_{j+1}=G_{j-1}$, it is immediate that $F^{\prime}_{j+2}=F_{j}$, $G^{\prime}_{j+2}=G_{j}$, and $P^{\prime}_{j+2}=P_{j}$, and the desired result follows. ∎
@@ 3.3 Impact
The above attack shows that, given just one PES-PCBC-encrypted message and knowledge of only a single plaintext block for this encrypted message, a ‘fake’ message can be constructed that will be guaranteed to pass the integrity checks. This fact meant that it has been recognised since 1996/97 that PES-PCBC should not be used.
@@ 4.3 Cryptanalysis
Lemma 1 suggests a way of forging an IOBC-encrypted message so that the final block will contain the correct MDC. However, the problem remains of discovering $g(G^{\prime}_{u-1})\oplus g(G_{v-1})$, as this is used in constructing the forged ciphertext in the statement of the lemma. Recacha [10] discusses this very point, and explains that making this difficult motivates the inclusion of the function $g$ in the design of IOBC — that is, if $g$ was not included (as is the case for PES-PCBC), then simple forgeries could be achieved.
@@
It is not hard to see that if $g^{k}(G_{j})=G_{j}$ for some $k$, then Lemma 2(ii) could be combined with Lemma 1 to yield a forgery attack (given a ciphertext message with corresponding known plaintext). This point is made by Recacha [10], who explains that the bit permutation $g$ has been chosen so that the smallest integer $i>1$ such that $g^{i}$ is the identity permutation is $(n/2-1)(n/2+1)=n^{2}/4-1$ (since $m=n/2$ is even). The restriction that the maximum length of messages encrypted using IOBC is $n^{2}/2-1$, as defined in Section 4.2, prevents this problem arising in practice.
@@
The above result can now be used in a straightforward way to enable message forgeries. As described in detail in [9], if $n=64$, given an IOBC ciphertext containing at least 685 blocks and some of the corresponding plaintext, then it is possible to create a forged ciphertext that will pass integrity checks with probability $2^{-22}$. 685 is clearly much less than the defined message length limit of 2047 blocks. A precisely analogous attack works for $n=128$, although the success probability is only $2^{-42}$.
@@ 5.2 A flawed cryptanalysis
The above Lemma implies that output bit pairs $(\ell^{\prime}_{i},r^{\prime}_{i})$ can never be equal to (0,1). In fact, we can obtain the following more general result which gives Lemma 4 as a special case.
Lemma 5 (Mitchell, [8])
@@
However, more recently, Di et al. [2] pointed out that the above analysis has a major flaw. The issue arises in the argument that, since $G_{j+1}=P_{j+1}\oplus F_{j}$, information about forbidden bit pairs in $F_{j}$, combined with knowledge of $P_{j+1}$, gives information about forbidden bit pairs in $G_{j+1}$. Di et al. [2] point out that if there are two possibilities for a bit pair in $F_{j}$ then there will always still be two possibilities for the corresponding bit pair in $g(G_{j+1})$ — as opposed to the analysis in [8] which suggests that the number of possibilities will be reduced to one with probability 1/6. That is, the number of possibilities for a bit pair in $g(G_{j+1})$ will never go below two, preventing the attack strategy working.
@@ 6.1 A chosen plaintext forgery attack
Under reasonable assumptions about the random behaviour of the encryption function, this will occur with probability $2^{-n}$. Hence, as $\Delta$ ranges over its $2^{n}$ possible values, the expected number of times that $C^{\prime}_{j}=C_{j}$ will hold is approximately 2, one of which will occur when $F^{\prime}_{j}=F_{i}$. The result follows. ∎
We can now give the following simple result that uses the same notation as Lemma 6. Note that it uses Lemma 1, which we already observed holds regardless of the choice of $g$.
Lemma 7
Suppose that $C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{t^{\prime}}$ and $C_{1},C_{2},\ldots,C_{t}$ are as in the statement of Lemma 6, and suppose also that $P^{\prime}_{j}=P_{i}$ and $C^{\prime}_{j}=C_{i}$ for some $j<t^{\prime}$ and $i<t$. Then, with probability approximately 0.5, the constructed ciphertext message
@@
The result follows immediately from Lemma 1, putting $u=j$, $v=i$ and observing that:
@@
which equals $C_{i}$ with probability approximately 0.5, since, by Lemma 6 we know that $G^{\prime}_{j-1}=G_{i-1}$ with probability approximately 0.5. ∎
@@ 6.2 A new attack approach with implications for EPBC
Suppose the ciphertext $C_{1},C_{2},\ldots,C_{t}$ was constructed using a scheme of the type shown in Figure 2, and that $P_{j}$ is a plaintext block for some $j$ satisfying $1<j<t$. Suppose the $(t+2)$-block ciphertext $C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{t+2}$ is constructed as follows:
@@
although the precise value of $P^{\prime}_{j+1}$ is unimportant. The result follows trivially. ∎
@@
Lemma 8
Suppose $g$ is as defined for EPBC, and (using the notation of Lemma 4) suppose also that $g(L||R)=L^{\prime}||R^{\prime}$. Then:
@@
Claim (i) follows immediately, and claim (ii) follows from observing that $\lnot\ell\land r=1$ if and only if $r=1$ and $\ell=0$. ∎
From Theorem 6.1, we can construct a possible forgery by guessing the value of $G_{j}\oplus g(G_{j})$. If $g$ is as defined for EPBC, then, from Lemma 8(i), we simply need to guess the first $n/2$ bits of $G_{j}\oplus g(G_{j})$. Moreover, if we restrict our guesses for this ‘first half’ to $n/2$-bit strings containing at most $n/8$ ones (where we assume that $n$ is a multiple of 8), then from Lemma 8(ii) we will have a better than evens chance of making a correct guess. The number of such strings is simply:
@@ 6.3 Issues with the use of Initialisation Vectors (IVs)
We can state the following result, which is essentially a special case of Lemma 1, and applies to all modes adhering to the design of Figure 2 and for which $g$ is linear (as is the case for IOBC and, trivially, PES-PCBC). In such a case, as observed in [9], the distributivity property $g(X\oplus Y)=g(X)\oplus g(Y)$ for any $X$ and $Y$ holds.
Lemma 10
@@
The result follows from Lemma 1, setting $u=v=3$. ∎
Thus if two ciphertexts are encrypted using the same IV, and a single plaintext block is known for each message, then a forgery can be constructed. Observe that this attack can be extended using Lemma 2(ii).
@@ 7 Other related modes
There was a gap of some 16 years before the first of these two additional schemes was made public — IOC (short for Input and Output Chaining) was made public by Recacha in 2013 [11]. IOC is clearly closely related to IOBC and EPBC, but was designed to avoid the known issues with these schemes. IOC was made public at a time when there was a renewed interest in the area, at least partly prompted by a NIST initiative on lightweight cryptography [^2] As discussed at https://csrc.nist.gov/Projects/Lightweight-Cryptography, NIST began investigating cryptography for constrained environments in 2013, and one of the goals was to find lightweight methods for authenticated encryption.. A slightly revised version was made public in early 2014, [14]. Cryptanalyses of both versions of the scheme first appeared in 2014 [1].